OpenStudy (anonymous):

Which of the following shows 8x3y + x2 - 14z - 2 + 5y2x written in standard form? -14z + 8x3y + 5y2x - 2 + x2 8x3y + 5y2x + x2 - 14z - 2 8x3y + x2 + 5y2x - 14z - 2 x2 - 2 - 14z + 5y2x + 8x2y @Directrix @ganeshie8

OpenStudy (anonymous):

We write it in the highest exponent to the lowest. 8x^3y = 4 x^2 = 2 -14z = 1 -2 = 1 5y^2x = 3

OpenStudy (anonymous):

So what do you get?

OpenStudy (anonymous):

B

Directrix (directrix):

Why B) 8x3y + 5y2x + x2 - 14z - 2 and not C) 8x3y + x2 + 5y2x - 14z - 2 @knightmare6

OpenStudy (anonymous):

its not right bc 5y2x is bigger than x2

Directrix (directrix):

The degree, yes. I agree with option B, then.

Directrix (directrix):

@knightmare6